OpenStudy (rockinhood):

1 2 ____ = ___ c+5 3

OpenStudy (anonymous):

are you familiar with proportions ?

OpenStudy (rockinhood):

Mostly, yes.

OpenStudy (anonymous):

what you have here is a proportion, and to solve this, you must cross multiply. 1/(c + 5) = 2/3 cross multiply 2(c + 5) = (1)(3) can you take it from here and solve for c. Don't forget to distribute through the parenthesis.

OpenStudy (anonymous):

2c + 10 = 3 2c = 3 - 10 2c = -7 c = -7/2 check. 1/(c + 5) = 2/3 1/(-7/2 + 5) = 2/3 1/(-7/2 + 10/2) = 2/3 1/(3/2) = 2/3 1 * 2/3 = 2/3 2/3 = 2/3 (correct) c = -7/2 any questions ?

OpenStudy (rockinhood):

I understand. Thanks a ton :D